# Decoding Quantum LDPC Codes using Collaborative Check Node Removal

### v1

@@ IV-A Weak check node removal
From here on we address the decoding algorithm for the CTS. Later, we show generalization to QTS requires modification in the removal rate of check nodes. We saw before that the CTS for the GHP code is rooted in the contribution of a single circulant matrix. For example, the CTS shown in Fig 1(a) is due to the circulant matrix $bI_{7}$. These types of trapping sets have odd-degree check nodes adjacent to them, which we exploit to initialize the decoding. Therefore, we identify the decoding problem as syndrome based and extend the definition of CT construction with unsatisfied check nodes as the roots. Further, we adopt the notion of week check nodes from [13], which assumes that at any level of $T_{\mathcal{K}}(c),\forall c\in\mathtt{UNSAT}$; there exist at least one weak check node, which is defined in IV.1.
Definition IV.1 (Weak check nodes).
@@
If in level $t$ of the computation tree $T_{\mathcal{K}}(c)$ has a root unsatisfied check node $c\in UNSAT$, then there exists at-least a check node $z$, which has the shortest path to any of the nearest qubit $v\in\mathcal{T}_{s}$, then $z$ is a weak check node at level $\mathcal{K}=t$ of $T_{\mathcal{K}}(c)$.
@@
Therefore, instead of removing the exact mis-satisfied check nodes from $T_{\mathcal{K}}(v)$, we remove the weak check nodes from level $\mathcal{K}=1$ of $T_{\mathcal{K}}(c_{i})$. It also improves the separation of the adjacent trapped qubit of the unsatisfied check node $c_{i}$. For example, Fig 5 in Appendix VI-A shows $c_{5}$ as the weak check node at $\mathcal{K}=1$. Even if the knowledge of $(3,3)$ CTS is unknown, simply removing $c_{5}$ from $\mathcal{K}=1$ of $T_{k}(c_{0})$ can increase the separation of the adjacent qubit $v_{0}$. This helps introduce a lot of freedom and simplicity to the decoding.
@@
Obviously one can remove weak check nodes from any level $\mathcal{K}$ of $T_{\mathcal{K}}(c_{i})\,\,\forall c_{i}\in\mathtt{UNSAT}$. Also it is obvious that not all the check nodes at level $t$ are the weak check nodes and removing these will have a negative affect in the further iterative decoding. Therefore, we use the parameter “deselection degree” denoted as $\mathtt{df}$, to restrict the number of random check node removal from any level of the CT. Also, note that alternatively one can identify the level $\mathcal{K}=t$ of the computation tree as the level of all leaf check nodes of the computation tree $T_{\mathcal{K}}(c_{i})$. This method of removing weak leaf nodes randomly is the key assistive part of the overall decoding architecture. The complete algorithm is known as the Fixed Node Check Removal (FNCR) [13], which is described in Algorithm 1.
@@
1 $\mathtt{rem}$ = [];
2for $\mathtt{check}$ in $\mathtt{UNSAT}$ do
@@ 3       $\mathtt{leaf}$leftarrow All the leaf nodes of $T_{t}(\mathtt{check})$;
4       $\mathtt{dis}$leftarrow random.chose($\mathtt{leaf}$,$\mathtt{df}$);
5       $\mathtt{rem}$circ $\mathtt{dis}$;
6foreach $\mathtt{uc}$in $\mathtt{rem}$ do
@@ 5       $\mathtt{rem}$circ $\mathtt{dis}$;
        // Row removal
@@
Algorithm 1 FNCR decoding
IV-B Collaborative architecture
For CTS, it is enough to set $\mathtt{df}=1$, as can be seen from Fig. 5 and the numerical results of appendix VI-A. For improving the qubit separation in QTS, ideally, the removal rate should be set to at least $\mathtt{df}=2$, as seen from Fig. 6 in appendix. The modified check matrix output of algorithm 1 has a high probability of having qubits with increased separation and thus offers better convergence with iterative decoders. Therefore, an iterative decoder can be used twice, first with the full parity check matrix and again on the modified parity check matrix from Algorithm 1. We call this approach a collaborative decoding, details of which are described in Algorithm 2. Collaborative decoding can be considered primarily as using a BP decoder in two modes. During the main decoding mode, BP uses the complete parity check matrix of the QLDPC code and the other mode is where it uses the modified parity check matrix output of Algorithm 1, for predicting an error estimate. The later is considered as the sub-decoding mode.
@@ IV-B Collaborative architecture
  • Maximum iterations of primary BP:= $\mathtt{max_{iter}}$,
  • Maximum iteraions of sub BP:= $\mathtt{max_{sub}}$,
@@
  • Sampling degree: $\mathtt{sf}$,
  • Maximum FNCR rounds: $\mathtt{fr}$,
@@
output : Total predicted error pattern $\mathtt{\hat{e}}$.
$\mathtt{unchanged=itn=0}$;
  // Main decoder init
@@ 1 $\mathtt{s_{prev}:=s}$;
3       $\mathtt{\hat{e}_{bp}}$ = $\mathsf{BP}$($\mathbf{H},\mathtt{llr},\mathtt{max_{iter}}$).$\mathsf{decode}$($\mathtt{s}$);
@@ 4       $\mathtt{\hat{s}_{bp}}$ = $\mathtt{\hat{e}_{bp}}\mathbf{H}^{T}$;
5       $\mathtt{itn}\leftarrow\mathtt{itn}+\mathtt{max_{iter}}$;
6       if $\mathtt{\hat{s}_{bp}}=\mathtt{s_{prev}}$ then
@@ 10      $\mathtt{s_{prev}\leftarrow\hat{s}_{bp}}$;
11
[…]
13 $\hat{e}_{sub}:=\mathsf{zeros}\mathtt{((}\mathtt{fr},\mathbf{H}.\mathsf{shape}\mathtt{[1]}\mathtt{))}$ ;
@@ 16       $\mathtt{s_{fn}\leftarrow s+\hat{s}_{bp}+(\sum_{j}s_{sub}[j]),\forall j<i}$;
17       if $\mathtt{check=0,\forall check\in s_{fn}}$ then
18             return $\mathtt{\hat{e}\leftarrow\hat{e}_{bp}+\sum_{i}\hat{e}_{sub}[i]}$;
@@ 19             End;
20
21      for $\mathtt{check}$ $\mathsf{in}$ $\mathtt{s_{fn}}$ do
22             if $\mathtt{check\neq 0}$ then
23                   $\mathtt{UNSAT}$.$\mathsf{extend}$($\mathtt{check}$);
24
25
26      $\mathtt{u_{s}=}\mathsf{random.sample}\mathtt{(UNSAT,sf)}$;
27       $\mathbf{H}_{FN}=\mathsf{FNCR}(\mathbf{H},\mathtt{u_{s}},\mathtt{t},\mathtt{df})$;
29       $\mathtt{\hat{e}_{sub}[i]=}\mathsf{BP_{sub}.decode}\mathtt{(s_{fn})}$;
30       $\mathtt{s_{sub}[i]\leftarrow\hat{e}_{sub}[i]}\mathbf{H}^{T}$;
@@ 31       $\mathtt{i}\leftarrow\mathtt{i}+\mathtt{1}$
return $\mathtt{\hat{e}\leftarrow\hat{e}_{bp}+\sum_{i}\hat{e}_{sub}[i]}$
Algorithm 2 Collaborative BP decoding with FNCR
@@
In Fig 3, we show the advantage of using BP in the sub-decoding mode. We observe, out of numerous sampled instances, a case where the main BP decoder gets stuck while decoding the syndrome of an error pattern generated at an independent bit flip physical error rate of $p=0.03$ for the $[[882,24]]$ GHP code after a few iterations. The stuck scenario represents the fact that the main decoder cannot identify anymore correct stabilizer violations. The tail of the plot shows that by activating the sub-decoder, we can predict the syndrome correctly and thus, eventually all the errors are correctly predicted. Therefore, we confirm that removing weak check nodes using Algorithm 1 increases the efficacy of the decoding output.
@@
Figure 3: Numerical experiment showing the sub-decoder is able to identify new correct unsatisfied stabilizers, which the main decoder alone can not predict. We observe the decoding performance for the syndrome obtained at independent bit flip physical error rate $p=0.03$ for $[[882,24]]$ GHP code. The main BP decoder can not predict any more correct stabilizer violations and gets stuck at $40^{th}$ iteration. Subsequently, through sub-decoding, we call Algorithm 1 repeatedly and BP is applied on the output of the same. We observe repeated calls of Algorithm 1 leads to an improvement to the performance of BP in the sub-decoding mode. We also observe a tail in the plot signifying newly accurate predictions of the unsatisfied stabilizer checks.
Further, we perform some memory experiments to compare the performance of Algorithm 2 with standard BP min-sum decoder and the state-of-the-art BP+OSD decoder. In Fig 4, we show the results, which indicate that the proposed decoding scheme offers significant improvements in the logical error rates over the BP decoder, indicating a potential breakthrough to circumvent the harmful configurations for the iterative BP. Algorithm 2 give logical error rates better than BP but it lags compared to BP+OSD. While BP+OSD uses an exhaustive search strategy to find the most accurate error patterns, Algorithm 2 is devoid of such high-order post-processing steps. There is still a scope to improve accuracy to identify check nodes. For any improvement, we note that the parameter $\mathtt{df}$ can help improve the decoding success rates.
Figure 4: The logical error rates obtained for the GHP codes from [6], under the independent bit flip Pauli channel. We perform the code capacity LER simulations for both cases. In (a) we benchmark with the $[[882,24]]$ GHP code and in (b) we benchmark the $[[1270,28]]$ GHP code. The LERs are plotted within a $95\%$ confidence intervals, which are indicated through the error-bars. In both cases, LERs are improved using Algorithm 2 over iterative min-sum BP.
IV-C Time Complexity
The collaborative decoding essentially has two phases, one consisting of the action of BP and the other consisting the process of modifying the parity check matrix for the sub-decoder. The cost of the latter mainly comes from the process of constructing the computation tree from the unsatisfied check nodes. Now following the concept of weak check nodes, we discussed that a computation tree of level $\mathcal{K}=1$ is enough to asses the possibility of increasing the separation. For a bi-regular $(d_{v},d_{c})$ QLDPC code, such a tree construction follows from adding the $dc$ adjacent data qubits to the root, which is a $\mathcal{O}(n)$ process. Further the leaf check nodes of level $\mathcal{K}=1$, are added in $\mathcal{O}(m\,d_{c})$ time, where $m$ is the number of stabilizers of the code. As QLDPC codes show extreme sparsity, i.e. $d_{c}<<<m$, Algorithm 1 or the process of generating the modified check matrix is linear, i.e. $\mathcal{O}(n)$ for a single unsatisfied stabilizer or check node. If the support of the syndrome is $\mathrm{supp}(s)=||s||$, then the algorithm has an overall time complexity of approximately $\mathcal{O}(||s||n)$. We also point out that the complexity of Algorithm 1 can even be improved through sampling the unsatisfied check nodes from the syndrome, which essentially reduces the cardinality of the input $\mathtt{UNSAT}$ for Algorithm 1. We use the parameter $\mathtt{sf}$, to determine the sampling rate of such unsatisfied check inputs.
The collaborative decoder can offer an almost linear computational complexity considering the syndrome weight is low (or, low physical error rate of qubits) or, a heavy sampling of the unsatisfied check nodes. Therefore, the time complexity of the overall collaborative decoding has an implicit dependence on the support of the received syndrome.

### v2

@@ IV-A Weak stabilizer check node removal
Here we mainly address a technique relevant to improving min-sum decoder’s performance in the context of CTS. Later, we show that the outline of weak stabilizer check node removal can smoothly merge in the overall decoding algorithm, when we also address the QTS. We saw before that the CTS for the GHP code is rooted in the contributions from a single circulant matrix. For example, the CTS shown in Fig 1(a) is due to the circulant matrix $bI_{7}$. These types of trapping sets have odd-degree check nodes adjacent to them, which we exploit to initialize the decoding. Therefore, the decoding problem is essentially a syndrome-based algorithm, which leverages the CT constructed with unsatisfied check nodes as the roots. We now adopt the notion of week check nodes from [16], which we will show later to be invaluable in the overall decoding architecture. It is defined as:
Definition IV.1 (Weak stabilizer check nodes).
@@
In the computation tree $T_{\mathcal{K}}(c_{r})$ with an unsatisfied check node $c_{r}\in\mathtt{UNSAT}$, as it’s root, has at-least one check node $c$ at any level $t$ of $T_{\mathcal{K}}(c_{r})$, which has the shortest path to any of the nearest qubit $v\in\mathcal{T}_{s}$, then $c$ is a weak stabilizer check node at level $\mathcal{K}=t$ of $T_{\mathcal{K}}(c_{r})$.
@@
Therefore, instead of removing the exact mis-satisfied check nodes from $T(v)$ for any $v\in\mathcal{T}_{s}$, we remove the weak stabilizer check nodes from level $\mathcal{K}=1$ of $T_{\mathcal{K}}(c_{r})\,\forall c_{r}\in\mathtt{UNSAT}$. This improves the separation of the adjacent trapped qubit of the unsatisfied check node $c_{r}$. For example, we observe in Fig 3 that $c_{5}$ is a weak stabilizer check node at $\mathcal{K}=1$ of $T(c_{0})$ for the $(3,3)$ CTS of the $[[882,24]]$ GHP code, shown in Fig. 1(a). This enable us to probabilistically sample the check nodes from level $\mathcal{K}=1$ of $T_{1}(c_{0})$ and a hit on $c_{5}$ increase the separation of the adjacent erroneous qubit $v_{0}$, which in-turn enables the min-sum decoder to predict the same, even if the knowledge of $(3,3)$ CTS is unknown. This introduce a lot of freedom and simplicity to the decoding.
@@
Obviously one can remove weak check nodes from any level of CT $T_{\mathcal{K}}(c_{i})\,\,\forall c_{i}\in\mathtt{UNSAT}$. We chose level $1$ of the CT as the only choice for removing check nodes. This choice corresponds to a universal level in the CT, where removing leaf check nodes serves the interest of both the CTS and QTS. Further, it is obvious that not all the check nodes at any level are the weak nodes and removing these does not have any positive effect on the min-sum decoder. We use the parameter “deselection degree” denoted as $\mathtt{df}$, to sample random check nodes from level $1$ of the CT and later remove them in the process. This method of removing weak leaf nodes randomly is the key assistive part of the overall decoding architecture, which enables the mitigation of harmful configurations without knowing the particulars of the underlying TS. The complete algorithm has a minor modification compared to it’s classical counterpart known as the Fixed Node Check Removal (FNCR) [16]. We call the modified algorithm simply the check node removal (CNR) algorithm, which is described in Algorithm 1.
@@
$\mathtt{rem}$ = [];
  // Initialize possible leaf checks to remove
1 for $\mathtt{check}$ in $\mathtt{UNSAT}$ do
@@ 2       $\mathtt{leaf}$leftarrow All the leaf nodes of $T_{t}(\mathtt{check})$;
3       $\mathtt{rem}$circ $\mathtt{leaf}$;
[…]
5 foreach $\mathtt{uc}$in $\mathtt{dis}$ do
@@ 3       $\mathtt{rem}$circ $\mathtt{leaf}$;
        // Row (check node) removal
6
return $\mathbf{H_{FN}}$ ;
  // The modified matrix
Algorithm 1 Check Node Removal (CNR)
IV-B Use of ‘Information Measurement’ for better efficacy
In the previous section, we discussed the CNR algorithm inspired by the classical probabilistic FNCR algorithm [16]. As one can observe, CNR acts on a large sample space of possible nodes to probabilistically select candidate check nodes for removal. Here the sample space refers to all the candidates, from which our algorithm randomly choses $\mathtt{df}$ number of stabilizer check nodes for removal. In algorithm 1, this space contains all the elements present in the $\mathcal{K}=1$ level of the CT of an unsatisfied check node. Although this method improves the iterative decoder’s performance within a finite number of iterations [16], for large codes, the number of calls to FNCR or CNR might increase. Further, we observe it can severely harm the success rate of the overall decoder. In this section we attempt to reduce the sample space of check nodes for probabilistic removal. The aim is to exclude any kind of check nodes from the sample space, which are incapable to commit any sort of improvement in the iterative decoder’s performance. This suggests a sample space, elements of which satisfy Lemma 2 regarding any QTS, and for CTS, these elements are the check nodes, which can correct any of the trapped data qubits of the underlying TS. We use the concept of information measurement (IM) to quantify the high value stabilizer check nodes at level $1$ of a CT of a root unsatisfied check node. We claim the nodes with the highest IM values are the check nodes, responsible for limiting the separation of erroneous trapped qubits, especially in the context of QTS. IM associates a metric to both the data qubits and stabilizer checks and is defined as follows:
Definition IV.2 (Information Measurement for data qubits).
For each data qubit of the quantum code, the information measurement is defined by the total number of adjacent stabilizer checks that are unsatisfied.
i.e.
$\mathrm{IM}_{v_{i}}=\sum_{j}c_{j},\forall c_{j}\in N(v_{i}),\text{if}\,\,c_{j}=1.$ (23)
Definition IV.3 (Information Measurement for Stabilizer checks).
For each stabilizer check, the information measurement is defined by the sum of information measurement values of all of its adjacent data qubits.
i.e.
$\mathrm{IM}_{c_{j}}=\sum_{i}\mathrm{IM}_{v_{i}},\forall v_{i}\in N(c_{j}).$ (24)
Consider the CT of $v_{351}$ shown in Fig. 4. Assume a typical harmful error configuration $e$ with support $\mathrm{supp}(e)=\{v_{0},v_{351},v_{405}\}$. Now, if we evaluate the IM values of the stabilizer check nodes at level $2$ of $T(v_{351})$, the $6$ check nodes shown in the figure are the ones with the highest IM values. Therefore, chosing the check nodes with maximum IM values eliminates the possibilities for any of the redundant check node’s presence during the random sampling. In appendix we provide algorithm 4, which describe the function ‘$\mathtt{FIND\_IMs}$’. This function returns the IM values of input stabilizer check nodes stored in a dictionary. We modify the algorithm 1 and propose Quantum Check Node Removal (QCNR) in algorithm 2.
input : $\mathbf{H}$, $\mathtt{UNSAT}$, $\mathtt{t}$, $\mathtt{df}$
output : Modified parity check matrix $\mathbf{H}_{FN}$
$\mathbf{H}_{FN}=\mathbf{H}$;
  // Initialization
$\mathtt{rem}$ = [];
  // Initialize checks removal
1 for $\mathtt{check}$ in $\mathtt{UNSAT}$ do
2       $\mathtt{leaf}$leftarrow All the leaf nodes of $T_{t}(\mathtt{check})$;
3       $\mathtt{leafval}\leftarrow\textnormal{{Find\_IMs}}(\mathtt{leaf})$;
       $\mathtt{maxleaf}=\mathrm{MAX}_{c}(\mathtt{leafval})$;
        // Check node with maximum IM value
4       $\mathtt{rem}$circ $\mathtt{maxleaf}$;
[…]
6 foreach $\mathtt{uc}$in $\mathtt{dis}$ do
        // check node removal
@@ 4       $\mathtt{rem}$circ $\mathtt{maxleaf}$;
Algorithm 2 Quantum Check Node Removal (QCNR)
This algorithm proceeds with selecting the keys with maximum value from the output of the algorithm ‘$\mathtt{FIND\_IMs}$’. These keys are the stabilizer check node indices with the highest IM values compared to all the leaf nodes of the CT. These selected check nodes then form the sample space of nodes, from which sampling of removable stabilizer check nodes is done according to the specified deselection degree, and ultimately, the quantum code’s Tanner graph is modified. Next, we describe the complete collaborative decoding architecture.
IV-C Collaborative architecture
We now leverage both the concepts of ‘weak stabilizer check nodes’ and ‘information measurement’ to rescue the min-sum decoder from the trapped scenarios. In previous sections, we discussed the improvement in the separation of trapped data qubits due to the modified Tanner graph obtained from CNR, in Algorithm 1. Then, we introduced QCNR, in Algorithm 2, which has an even better probability of having trapped qubits with increased separation, therefore offering better convergence with iterative decoders.
We now propose the use of QCNR in a collaborative setting to deduce a low complexity decoding architecture. We mainly use the iterative or, min-sum decoder in two modes. The main mode implements the min-sum BP on the unmodified parity check matrix of the QLDPC code. The other mode, which we call sub-decoding mode, first removes some stabilizer check nodes from the quantum code’s Tanner graph using Algorithm 2. The min-sum BP is then implemented on the resultant modified parity check matrix of the QLDPC code. The decoding process is initialized with the main decoding mode. As we discussed extensively in previous sections, the main decoding mode is limited in its performance. Therefore, if for a predetermined number of iterations of BP, no more correct predictions are done by the decoder, it switches to the sub-decoding mode. After the main and sub decoding rounds are completed, the decoder switches again to the main decoding mode. We note that in any round of the main and sub decoding rounds, we consider the corrections of all previous rounds, i.e. we continuously update the list of the net unsatisfied stabilizer check nodes. Therefore, calling the main decoding mode again after the completion of a sub-decoding mode is to correct any error that does not require any further sub-decoding (this is meant to tackle errors that are no longer supported over any TS after the previous main+sub-decoding round). This approach is different than the post-processing based methods, as it switches its message-passing operation over different parity check matrices and outputs a prediction only if all the errors are corrected or a predetermined number of sub-decoding rounds have been completed. We describe the complete details of this collaborative decoding in Algorithm 3.
@@ IV-C Collaborative architecture
  • Maximum iterations of main BP:= $\mathtt{max_{iter}}$,
  • Maximum iterations of sub BP:= $\mathtt{max_{sub}}$,
@@
  • Maximum sub-decoding rounds: $\mathtt{fr}$,
@@
output : Predicted error pattern $\mathtt{\hat{e}}$.
1 $\mathtt{unchanged=0}$;
$\mathtt{itn}=1$;
  // Main decoder initialization
@@ 2 $\mathtt{s_{prev}:=s}$;
4       $\mathtt{\hat{e}_{bp}}$ = $\mathsf{BP}$($\mathbf{H},\mathtt{llr},\mathtt{itn}$).$\mathsf{decode}$($\mathtt{s}$);
@@ 5       $\mathtt{\hat{s}_{bp}}$ = $\mathtt{\hat{e}_{bp}}\mathbf{H}^{T}$;
6       $\mathtt{itn}\leftarrow\mathtt{itn}+1$;
7       if $\mathrm{supp}(\mathtt{\hat{s}_{bp}})=\mathrm{sup}(\mathtt{s_{prev}})$ then
@@ 11      $\mathtt{s_{prev}\leftarrow\hat{s}_{bp}}$;
12
$\mathtt{e_{net}}:=\mathsf{zeros}\mathtt{((}\mathtt{fr},\mathbf{H}.\mathsf{shape}\mathtt{[0]}\mathtt{))}$;
  // Sub-decoder init
13 $\mathtt{s_{net}}:=\mathsf{zeros}\mathtt{((}\mathtt{fr},\mathbf{H}.\mathsf{shape}\mathtt{[0]}\mathtt{))}$;
@@ 16       $\mathtt{s_{fn}\leftarrow s+\hat{s}_{bp}+(\sum_{j}s_{sub}[j]),\forall j<i}$;
17       $\mathtt{UNSAT}$.$\mathsf{extend}$($\mathrm{supp}(s_{fn})$);
18
19      $\mathbf{H}_{FN}=\mathsf{QCCNR}(\mathbf{H},\mathtt{UNSAT},\mathtt{t},\mathtt{df})$;
21       $\mathtt{\hat{e}_{sub}=}\mathsf{BP_{sub}.decode}\mathtt{(s_{fn})}$;
22       $\mathtt{s_{sub}\leftarrow\hat{e}_{sub}}\mathbf{H}^{T}$;
23       $\mathtt{\hat{e}^{\prime}_{bp}}$ = $\mathsf{BP}$($\mathbf{H},\mathtt{llr},\mathtt{max_{iter}}$).$\mathsf{decode}$($\mathtt{s_{sub}+s_{fn}}$);
24       $\mathtt{\hat{s}^{\prime}_{bp}}$ = $\mathtt{\hat{e}^{\prime}_{bp}}\mathbf{H}^{T}$;
25       $\mathtt{e_{net}}[i]\leftarrow\mathtt{e_{sub}}+\mathtt{\hat{e}^{\prime}_{bp}}$;
26       $\mathtt{s_{net}}[i]\leftarrow\mathtt{s_{sub}}+\mathtt{\hat{s}^{\prime}_{bp}}$;
27       if $\mathrm{supp}(\mathtt{s_{net}}[i])=\emptyset$ then
28             return $\mathtt{\hat{e}\leftarrow\hat{e}_{bp}+\sum_{m:m\leq i}\mathtt{e_{net}}[m]}$;
@@ 29             End;
30
@@ 31      $\mathtt{i}\leftarrow\mathtt{i}+\mathtt{1}$
return $\mathtt{\hat{e}\leftarrow\hat{e}_{bp}+\sum_{i}\hat{e}_{net}[i]}$
Algorithm 3 Quantum Collaborative Check Node Removal (QCCNR) Algorithm
@@
In Fig. 5, we show the advantage of using BP in the sub-decoding mode. We observe, out of numerous sampled instances, a case where the main BP decoder gets stuck while decoding the syndrome of an error pattern generated at an independent bit flip physical error rate of $p=0.03$ for the $[[882,24]]$ GHP code after a few iterations. We note that for all the numerical experiments of GHP codes, we use the following parameter values: maximum-BP-iterations = $100$, BP-method = “minimum-sum”, scaling-factor = $0.625$. These parameters are relevant to the message passing of the min-sum BP in both the main and sub-decoding modes. The stuck scenario represents the fact that the main decoder cannot identify any more correct stabilizer check violations. After the main mode of BP decoder fails to correct any more errors for a consecutive $11$ rounds, we iteratively apply Algorithm 2 and run the BP decoder on a modified parity check matrix. We keep track of the new corrections at each round of the sub-decoding. The tail of the plot indicates that by activating the sub-decoding mode, new stabilizer violations are predicted correctly and thus, eventually, all the errors are corrected. This numerically confirms that removing stabilizer check nodes using Algorithm 2 increases the efficacy of the decoding output.
@@
Figure 5: Numerical experiment showing the sub-decoder’s ability to identify stabilizer violations correctly, which were not possible by decoding with min-sum based BP alone (which we denote as ‘main decoding region’). We assume that the error occurs on the data qubits after each round of an error correction cycle, and the syndrome measurements are perfect. Under these phenomenological noise assumptions, we sample an error pattern at an independent bit flip physical error rate of $p=0.03$ for the $[[882,24]]$ GHP code. The min-sum BP decoder can not predict any more correct stabilizer violations for a consecutive $11$ iterations and gets stuck at the $40^{th}$ decoding iteration. Subsequently, we initialize the sub-decoding mode. We call Algorithm 2 repeatedly, and min-sum BP is applied on the output parity check matrix of the QCNR algorithm. To tackle all the trapped cases for QTS and CTS, we first set the deselection degree $\mathtt{df}=6$ and then set $\mathtt{df}=1$ for the last 20 sub-decoding iterations. We observe repeated calls of Algorithm 2 lead to an improvement in the performance of BP in the sub-decoding region. The tail in the plot signifies newly accurate predictions of the violated stabilizer checks and therefore indicates the improved performance.
Further, we perform memory experiments to compare the performance of the proposed QCCNR decoder in Algorithm 3 with the min-sum BP decoder and the state-of-the-art BP+OSD decoder with the order of OSD set to 00. For the decoding of GHP codes in the memory experiments, we used a strategy where we set the maximum number of sub-decoding rounds to $200$. This implicitly includes the case where, if there are no more errors to correct, the decoder stops and outputs the prediction, which is also indicated in Algorithm 3. Now, we have adopted a scenario where in the first $100$ sub-decoding rounds we set the deselection degree to $\mathtt{df}=6$ and for the rest of $100$ rounds we set $\mathtt{df}=1$. This strategy is adopted deliberately to first target the errors supported on any QTS and then to target the errors supported on the CTS. The degree of the data qubits is $d_{v}=3$, for the family of GHP codes we chose. Therefore, from Lemma 2, it is clear that to have the best possible result, we need to set the deselection degree to $\mathtt{df}=d_{v}(d_{v}-1)=6$. In Fig 6, we show the results, which indicate that the proposed decoding scheme offers significant improvements in the logical error rates over the min-sum BP decoder, indicating a potential breakthrough to circumvent the harmful configurations for the iterative min-sum BP. Algorithm 3 gives logical error rates much better than standard min-sum BP and almost achieves a performance comparable to the OSD00 post-processor.
Figure 6: The logical error rates obtained for the GHP codes from [7], under the independent Pauli-$\mathrm{X}$ noise channel. We perform the code capacity LER simulations for both cases. In (a) we benchmark with the $[[882,24]]$ GHP code and in (b) we benchmark the $[[1270,28]]$ GHP code. In both the experiments we use the min-sum algorithm of BP and for benchmarking against the OSD decoder, we use the the $0^{\text{th}}$ order OSD, i.e. BP+OSD00. For the QCCNR algorithm, we use the following parameters for Algorithm 3: $\mathtt{max_{iter}}=\mathtt{max_{sub}}=100,\mathtt{fr}=200$. Also, for the first $100$ sub-decoding rounds, the deselection degree is set as $\mathtt{df}=6$. This is set from Lemma 3 to support qubits trapped inside QTS. Further, for the remaining $100$ sub-decoding rounds, the deselection degree is set to $\mathtt{df}=1$, in support of the qubits trapped inside CTS, which follows from the section IV-A. We observe that the proposed QCCNR algorithm provides OSD-like improvements and obviously is far superior to the standard min-sum algorithm-based BP decoder.
IV-D Time Complexity
Algorithm 3 has two modes, as discussed previously. In both main and sub-decoding mode, the message passing of the BP decoder is linear, i.e., a $\mathcal{O}(n)$ process. In sub-decoding mode, Algorithm 2 has two parts: one is to construct the computation tree for each of the unsatisfied check nodes, and the other is to calculate the IM values and store them in a dictionary. The IM values evaluated for a particular node are an $\mathcal{O}(n)$ process and can be stored once. Essentially, for each round of the sub-decoding, we only require the IM values for a constant number of leaf check nodes, in the CT of each of the remaining unsatisfied check nodes. Therefore, the whole IM value evaluation takes at most $\mathcal{O}(|\mathtt{UNSAT}|n)$ amount of time. For GHP codes this time amounts to $\mathcal{O}(\frac{n^{2}}{\log{n}})$ [2]. The other process involved in Algorithm 2 is the construction of a level $1$ computation tree for each of the unsatisfied check nodes. If the QLDPC code is $(d_{v},d_{c})$ regular, constructing a level $1$ computation tree from any particular root check node requires first adding the $d_{c}$ number of data qubits and then for each of those data qubits adding $d_{v}$ number of stabilizer check nodes. This whole process is a $\mathcal{O}(n)$ process, as $d_{v},d_{c}<<n$ and are constants. Therefore, the overall cost of the decoder is dependent on the maximum number of unsatisfied stabilizers as indicated by the syndrome. For GHP codes, we approximate that this will result in a sub-quadratic decoder. In any case, if $|\mathtt{UNSAT}|$ grows linearly, then the decoder is quadratic. Therefore, QCCNR provides an excellent strategy for improving the iterative decoder’s performance compared to decoders like BP+OSD0 or the higher order alternatives.